# Odd Order Genuinely Positive Semi-Definite Tensors and Strongly Positive Semi-Definite Tensors

### v1

@@ Abstract
An even order symmetric tensor is positive semi-definite (PSD) if and only if it has no negative H-eigenvalues. As an extension, a tensor is said to be a generalized PSD tensor if its corresponding symmetrization tensor has no negative H-eigenvalue. In the odd order case, this extension contains a peculiar tensor class, whose members have no H-eigenvalues at all. We call such tensors barren tensors, and the other generalized PSD symmetric tensors genuinely PSD symmetric tensors, which contain useful structured tensors, such as symmetric M-tensors, completely positive tensors and Laplacian tensors of uniform hypergraphs. However, there are exceptions. It is known an even order symmetric B-tensor is a PSD tensor. We give an example of an odd order symmetric B-tensor, which is a barren tensor. We further introduce strongly PSD tensors which map real vectors to nonnegative vectors, and show that a strongly PSD tensor is a genuinely PSD tensor. An odd order completely positive tensor is a strongly PSD tensor. The set of strongly PSD tensors form a closed convex cone. However, the Laplacian tensor of a uniform hypergraph may not be strongly PSD.
Key words. Odd order tensors, positive semi-definite, symmetric M-tensors, completely positive tensors, Laplacian tensors, B-tensors, H-eigenvalues.
@@ 1 Introduction
We have another way to consider positive semi-definiteness of odd-order tensors.
Definition 1.2.
@@
In Section 3, we show that strongly positive semi-definite tensors are genuinely positive semi-definite tensors, and present their properties. We also show that completely positive semi-definite tensors are strongly positive semi-definite tensors, a strongly positive semi-definite tensor may have negative entries, and the Laplacian tensor of an odd order uniform hypergraph may not be a strongly positive semi-definite tensor. The set of strongly positive semi-definite tensors of the same order and dimension form a closed convex cone. We also study its dual cone.
@@ 4 Final Remarks
In this paper, we introduced odd order genuinely positive semi-definite tensors, barren tensors and strongly positive semi-definite tensors, and studied their properties. Several outstanding questions remain. We list them here.
@@
3. Hankel tensors are important in applications [12]. Even order strong Hankel tensors and even order complete tensors are positive semi-definite [14]. How about odd order strong Hankel tensors and odd order complete Hankel tensors?
4. Besides completely positive tensors, are there any other meaningful classes of strongly positive semi-definite tensors?

### v2

# Strongly Positive Semi-Definite Tensors and Strongly SOS Tensors

@@ Abstract
We introduce odd-order strongly PSD (positive semi-definite) tensors which map real vectors to nonnegative vectors. We then introduce odd-order strongly SOS (sum-of-squares) tensors. A strongly SOS tensor maps real vectors to nonnegative vectors whose components are all SOS polynomials. Strongly SOS tensors are strongly PSD tensors. Odd order completely positive tensors are strongly SOS tensors. We also introduce strict Hankel tensors, which are also strongly SOS tensors. Odd order Hilbert tensors are strict Hankel tensors. However, the Laplacian tensor of a uniform hypergraph may not be strongly PSD. This motivates us to study wider PSD-like tensors. A cubic tensor is said to be a generalized PSD tensor if its corresponding symmetrization tensor has no negative H-eigenvalue. In the odd order case, this extension contains a peculiar tensor class, whose members have no H-eigenvalues at all. We call such tensors barren tensors, and the other generalized PSD symmetric tensors genuinely PSD symmetric tensors. In the odd order case, genuinely PSD tensors embrace various useful structured tensors, such as strongly PSD tensors, symmetric M-tensors and Laplacian tensors of uniform hypergraphs. However, there are exceptions. It is known an even order symmetric B-tensor is a PSD tensor. We give an example of an odd order symmetric B-tensor, which is a barren tensor.
Key words. Odd order tensors, positive semi-definite, sum of squares, strict Hankel tensors, symmetric M-tensors, completely positive tensors, Hilbert tensors.
AMS subject classifications. 47J10, 15A18, 47H07, 15A72.
@@ 1 Introduction
In this paper, we consider positive semi-definiteness of odd-order tensors, by introducing the following tensor class.
Definition 1.1.
We show that completely positive semi-definite tensors are strongly positive semi-definite tensors, a strongly positive semi-definite tensor may have negative entries, and the Laplacian tensor of an odd order uniform hypergraph may not be a strongly positive semi-definite tensor. The set of strongly positive semi-definite tensors of the same order and dimension form a closed convex cone. We also study its dual cone.
A property stronger than positive semi-definiteness is the SOS (sum-of-squares) property, which is somehow associated with David Hilbert [7]. Hence, for odd order cubic tensors, we further introduce strongly SOS tensors. Strongly SOS tensors are strongly positive semi-definite tensors. We show that completely positive tensors are strongly SOS tensors. We introduce strict Hankel tensors which are also strongly SOS tensors. We show that the Hilbert tensors are strict Hankel tensors.
As some important structured tensors, such as the Laplacian tensor of an odd order uniform hypergraph, may not be a strongly positive semi-definite tensors, we thus consider some wider positive semi-definiteness-like properties for odd order tensors.
@@
In Section 3, we show that strongly positive semi-definite tensors are genuinely positive semi-definite tensors, and present their properties.
We introduce strongly SOS tensors and study their properties in Section 4.
@@ 3 Strongly Positive Semi-Definite Tensors
4 Strongly SOS Tensors and Strict Hankel Tensors
We now give the following definition.
Definition 4.1.
Suppose that $m$ is odd. Let $\mathcal{A}\in{S_{m,n}}$. Define $F:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ as $F(\mathbf{x})=\mathcal{A}\mathbf{x}^{m-1}$ for $\mathbf{x}\in{\mathbb{R}}^{n}$. We say that $\mathcal{A}$ is a strongly SOS (sum-of-squares) tensor if $F_{i}(\mathbf{x})$ is an SOS polynomial for $i=1,\dots,n$.
By definition, a strongly SOS tensor is a strongly positive semi-definite tensor. Furthermore, we have the following proposition.
Proposition 4.2.
Suppose $\mathcal{A}\in S_{m,n}$ is a strongly positive semi-definite tensor. If either of the following three cases holds:
  • (i)
    $n=2$,
  • (ii)
    $m=3$,
  • (iii)
    $m=5$ and $n=3$,
then $\mathcal{A}$ is also a strongly SOS tensor.
Proof.
As pointed out by David Hilbert [7], $F_{i}(\mathbf{x})$ is an SOS polynomial in all the above three cases. ∎
Corollary 4.3.
Example 3.8 is also a strongly SOS tensor. Thus a strongly SOS tensor may contain negative entries.
We now consider completely positive tensors.
Theorem 4.4.
A completely positive tensor $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in S_{m,n}$ is a strongly SOS tensor.
Proof.
$A(i)=\sum_{l=1}^{r}u_{i}^{(l)}{\left(\mathbf{u}^{(l)}\right)^{(m-1)}.}$
Therefore, $F_{i}(\mathbf{x})=\sum_{l=1}^{r}u_{i}^{(l)}\left(\left(\left(\mathbf{u}^{(l)}\right)^{\top}\mathbf{x}\right)^{\frac{m-1}{2}}\right)^{2}$. This completes the proof. ∎
Hankel tensors are important in applications [13]. In particular, even order strong Hankel tensors and SOS tensors [15]. We now extend this to the odd order case.
A Hankel tensor $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in T_{m,n}$ is a symmetric tensor, with a generating vector ${\bf h}=\left(h_{0},\dots,h_{(n-1)m}\right)^{\top}\in{\mathbb{R}}^{(n-1)m+1}$, such that
$a_{i_{1}\dots i_{m}}=h_{i_{1}+\dots+i_{m}},$
for $i_{1},\dots,i_{m}=0,\dots,n-1$. Then ${\bf h}$ is also a generating vector of a Hankel matrix $A=(a_{ij})\in T_{2,\lceil{(n-1)m\over 2}\rceil+1}$, defined by
$a_{ij}=h_{i+j},$
for $i,j=0,\dots,\lceil{(n-1)m\over 2}\rceil$. Here, if $m$ is odd and $n$ is even, $h_{(n-1)m+1}$ is an arbitrary real number. We say that $A$ is the associated Hankel matrix of the Hankel tensor $\mathcal{A}$. If $A$ is positive semi-definite, then we say that $\mathcal{A}$ is a strong Hankel tensor. An even order strong Hankel tensor is an SOS tensor. See Theorem 5.55 of [15].
Now, let $\mathcal{A}=\left(a_{i_{1}\dots i_{m}}\right)\in S_{m,n}$ be an odd order Hankel tensor, i.e., $m$ is odd. Denote $\mathcal{A}_{i}=\left(a_{ii_{2}\dots i_{m}}\right)\in S_{m-1,n}$ be the $(m-1)$th order Hankel tensor, with the first index $i_{1}$ being fixed at $i_{1}=i$. If $\mathcal{A}_{i}$ for $i=0,\dots,m-1$, are all strong Hankel tensor, then $\mathcal{A}$ is called a strict Hankel tensor. Clearly, strict Hankel tensors are strongly SOS tensors. We have the following theorem to display some strict Hankel tensors.
Theorem 4.5.
An odd order Hilbert tensor $\mathcal{A}\in S_{m,n}$ is a strict Hankel tensor, thus a strongly SOS tensor.
Proof.
Consider the associated Hankel matrix $A_{i}$ for $\mathcal{A}_{i}$, with $i=0,\dots,n-1$. Then any principal minor of $A_{i}$ is a principal minor of a big Hilbert matrix. Such a principal minor must be positive as the Hilbert matrix is positive definite [15]. This implies that $A_{i}$ is positive definite. Then $\mathcal{A}_{i}$ is a strong Hankel tensor for $i=0,\dots,n-1$. Hence, $\mathcal{A}$ is a strict Hankel tensor, thus a strongly SOS tensor. ∎
We draw the relations among strongly positive semi-definite tensors, strongly sum of squares tensors, completely positive tensors, and strict Hankel tensors for odd order tensors in Fig. 2.
Figure 2: The relations among strongly positive semi-definite tensors, strongly sum of squares tensors, completely positive tensors, and strict Hankel tensors for odd order tensors.
@@ 5 Final Remarks
In this paper, we introduced strongly positive semi-definite tensors and strongly SOS tensors, and studied their properties. We also introduced genuinely positive semi-definite tensors and barren tensors, and studied their properties. Several outstanding questions remain. We list them here.
@@ 3. Is the dual cone of SPSDm,n equal to SVm,n and whether SPSD3,n is a self-dual cone?
4. Give an example of a strongly positive semi-definite tensor, which is not a strongly SOS tensor.